# Decentralized Federated Learning: Model Update Tracking Under Imperfect Information Sharing

### 20 Mar 2024 v1

@@ Abstract
A novel Decentralized Noisy Model Update Tracking Federated Learning algorithm (FedNMUT) is proposed, which is tailored to function efficiently in the presence of noisy communication channels that reflect imperfect information exchange. This algorithm uses gradient tracking to minimize the impact of data heterogeneity while minimizing communication overhead. The proposed algorithm incorporates noise into its parameters to mimic the conditions of noisy communication channels, thereby enabling consensus among clients through a communication graph topology in such challenging environments. FedNMUT prioritizes parameter sharing and noise incorporation to increase the resilience of decentralized learning systems against noisy communications. Through theoretical and empirical validation, it is demonstrated that the performance of FedNMUT is superior compared to the existing state-of-the-art methods and conventional parameter-mixing approaches in dealing with imperfect information sharing. This proves the capability of the proposed algorithm to counteract the negative effects of communication noise in a decentralized learning framework.
Index Terms:
Optimization algorithms, Distributed control, Stochastic systems, Numerical algorithms
@@ I Introduction
In contemporary real-world scenarios, vast amounts of data are generated from various sources such as computers, mobile devices, smartwatches, and vehicles. These data are aggregated in centralized data centers for the purpose of training machine learning models. However, this centralized approach encounters significant hurdles, including limited communication bandwidth and privacy concerns, rendering it unreliable and non-scalable. Consequently, there is a pressing need to establish learning paradigms that not only ensure data security but also uphold privacy standards. This necessity has spurred the development of decentralized optimization algorithms, exemplified by Decentralized Federated Learning (DFL) and Federated Learning (FL). These methodologies have been increasingly applied across diverse sectors, including smart cities and connected vehicles [1, 2, 3, 4, 5, 6, 7, 8, 9].
I-A Literature Overview
Decentralized optimization typically employs consensus-based gradient descent methods, where the computed parameters are shared among clients [10, 1, 11]. These clients independently calculate local weights and gradients based on their data, subsequently exchanging these parameters with others. The aggregation of these parameters, influenced by the network’s topology which governs the communication framework, is pivotal in this learning paradigm. This network topology is often depicted as a simple graph, illustrating the communication pathways among clients. Decentralized approaches aim to alleviate the limitations inherent in centralized systems, such as communication delays and bandwidth constraints, thereby enhancing scalability and efficiency in extensive networks.
@@ I-A Literature Overview
Our preceding work delineated three algorithms for DFL under imperfect communication conditions, typified by noisy channels. These algorithms—FedNDL1, FedNDL2, and FedNDL3—differ in their handling of noise and parameter sharing, showcasing varying degrees of resilience to communication noise [24].
This paper introduces a novel algorithm that employs the Gradient Tracking method in DFL, considering the impact of communication noise. Previous studies have assessed the efficacy of two-time scale methods in DFL with noisy channels, but these investigations were constrained by rigid assumptions that are seldom met in practical, large-scale learning scenarios [25, 26, 27, 28, 29].
I-B Paper’s Contributions
@@
We evaluate the performance of decentralized FL using the Federated Model Update Tracking algorithm, which incorporates noise into tracking parameter transmissions to clients or servers. Unlike prior models where communication noise was not considered, our approach integrates noise post-local SGD updates, facilitating parameter exchange via a gossip/mixing matrix and updating the global parameters thereafter. We also compare the noise resilience of our algorithm with previously developed algorithms, providing both theoretical and empirical evidence of its robustness against communication noise [24].We show that the proposed algorithm handles the communication noise well than the FedNDL3 algorithm.
II Problem Statement
In this section, we describe the problem structure, assumptions, and the proposed algorithms that we analyzed in this paper. We start with a standard DFL setup in which $n$ clients have their own local datasets and collaborate with each other to update the global parameters. Formally, the problem can be represented as
$\min_{x\in\mathbb{R}^{d}}\Big{[}f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x_{i})\Big{]},$ (1)
@@ II Problem Statement
where $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ for $i\in\{1,\dots,n\}$ is the local objective function of the $i^{th}$ client node. The stochastic formulation of the local objective function can be written
$f_{i}(x_{i})=\operatorname{\mathbb{E}}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x_{i},\xi_{i})],$ (2)
@@
where $\xi_{i}$ is the data that has been sampled from the data distribution $\mathcal{D}_{i}$ for the $i^{th}$ client. The function $\ell(x_{i},\xi_{i})$ is the loss function evaluated for each client and for each data sample $\xi_{i}$. Here $x_{i}\in\mathbb{R}^{d}$ is the parameter vector of client $i$, and $X\in\mathbb{R}^{d\times n}$ is the matrix formed using these parameter vectors. The primary objective of the clients is to achieve optimality through collaboration i.e., $x_{i}=x^{*}=\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}f(x)$, which is a global minimizer.
@@
The mixing/gossip matrix, $W=[w_{ij}]\in[0,1]^{n\times n}$, is a non-negative, symmetric $(W=W^{\top})$ and doubly stochastic $(W\mathds{1}=\mathds{1},\mathds{1}^{\top}W=\mathds{1}^{\top})$ matrix, where $\mathds{1}$ is the column vector of unit elements of size $n$
FedNDL1 [24]: In this algorithm, each client in parallel performs updates first and then communicates the updated parameters to their neighbors. The communication depends on the topology of the communication graph, i.e., the mixing matrix, $W$, through a noisy communication channel. Due to the noisy communication channel, the neighboring client receives a noisy version of the parameters,
@@
where $\delta_{j}^{(t)}\in\mathbb{R}^{d}$, is a zero mean random noise and $x_{j}^{(t+\frac{1}{2})}$ is the vector of parameters sent by client $j$. Since we assume the noise to have a zero mean, the noise variance is
$D^{2}_{t,j}=\operatorname{\mathbb{E}}[\|\delta_{j}^{(t)}\|^{2}].$ (4)
FedNDL3 [24]: In this algorithm, the clients share their gradients over a noisy communication channel instead of the weights followed by the SGD update. This idea comes from our Noisy-FL motivation and the fact that SGD is inherently a noisy process. So, pursuing this scenario gives more flexibility to handle the noise as a part of the SGD process.
@@
FedNMUT: We next describe our proposed algorithm studied in this paper.
$x_{i}^{(t+1)}=x_{i}^{(t)}-\eta_{t}\widetilde{y_{i}}^{t}\$
$x_{i}^{(t+1)}=x_{i}^{(t)}-\eta_{t}({y_{i}}^{t}+\delta_{i}^{t})\$ (6)
$y_{i}^{t}=\Delta_{i}^{t}+\mu\Big{[}\sum\limits_{j\in\mathcal{N}(i)}w_{ij}(\widetilde{y_{j}}^{t-1}-\frac{1}{\eta}(\hat{x}_{j}^{t}-x_{i}^{t}))-\Delta_{i}^{t-1}\Big{]}\$
$\Delta_{i}^{t}=g_{i}^{t}-\frac{1}{\eta}\sum_{j\in\mathcal{N}(i)}w_{ij}(\hat{x}_{j}^{t}-x_{i}^{t})$
In this algorithm, $\mu$ is FedNMUT scaling factor/hyper-parameter that needs to be tuned to provide optimal solution. As noticed in Figure  1 and Figure 2, the convergence of the algorithm is dependent on the value of $\mu$.
Algorithm FedNMUT - Noisy Model Update Tracking
@@
1:  Input: For each node $i$ initialize: $x_{i}^{(0)}\ \in\ \mathbb{R}^{d}$, step size $\{\eta_{t}\}_{t=0}^{T-1}$, neighbors’ copy $\hat{x}_{j}^{0}$, step size $\eta$, GUT scaling factor $\mu$, mixing matrix $W=[w_{ij}]_{i,j\in[1,n]}$, $\mathcal{N}(i)$ represents neighbors of $i$ including itself, noise from the communication channel $\delta^{(t)}$
@@
Figure 1: Loss vs. iterations without noise for various $\mu$ values for different communication topologies.

### 24 Mar 2024 v2

# FedNMUT—Federated Noisy Model Update Tracking Convergence Analysis

@@ Abstract
A novel Decentralized Noisy Model Update Tracking Federated Learning algorithm (FedNMUT) is proposed that is tailored to function efficiently in the presence of noisy communication channels that reflect imperfect information exchange. This algorithm uses gradient tracking to minimize the impact of data heterogeneity while minimizing communication overhead. The proposed algorithm incorporates noise into its parameters to mimic the conditions of noisy communication channels, thereby enabling consensus among clients through a communication graph topology in such challenging environments. FedNMUT prioritizes parameter sharing and noise incorporation to increase the resilience of decentralized learning systems against noisy communications. Theoretical results for the smooth non-convex objective function are provided by us, and it is shown that the $\epsilon-$stationary solution is achieved by our algorithm at the rate of $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$, where $T$ is the total number of communication rounds. Additionally, via empirical validation, we demonstrated that the performance of FedNMUT is superior to the existing state-of-the-art methods and conventional parameter-mixing approaches in dealing with imperfect information sharing. This proves the capability of the proposed algorithm to counteract the negative effects of communication noise in a decentralized learning framework.
@@ I Introduction
I-A Motivation
In the real world, vast amounts of data are generated from various sources such as computers, mobile devices, smartwatches, and vehicles. These data are aggregated in centralized data centers for the purpose of training machine learning models. However, this centralized approach encounters significant hurdles, including limited communication bandwidth and privacy concerns, rendering it unreliable and non-scalable. Consequently, there is a pressing need to establish learning paradigms that not only ensure data security but also uphold privacy standards. This necessity has spurred the development of decentralized optimization algorithms, exemplified by Decentralized Federated Learning (DFL) and Federated Learning (FL). These methodologies have been increasingly applied across diverse sectors, including smart cities and connected vehicles [1, 2, 3, 4, 5, 6, 7, 8, 9]. In this paper, we propose a novel Decentralized Noisy Model Update Tracking Federated Learning algorithm (FedNMUT) that is tailored to function efficiently in the presence of noisy communication channels that reflect imperfect information exchange.
I-B Literature Overview
Decentralized optimization typically employs consensus-based gradient descent methods, where the computed parameters are shared among clients [10, 1, 11, 12]. These clients independently calculate local weights and gradients based on their data, subsequently exchanging these parameters with others. The aggregation of these parameters, influenced by the network’s topology which governs the communication framework, is pivotal in this learning paradigm. This network topology is often depicted as a simple graph, illustrating the communication pathways among clients. Decentralized approaches aim to alleviate the limitations inherent in centralized systems, such as communication delays and bandwidth constraints, thereby enhancing scalability and efficiency in extensive networks.
@@ I-B Literature Overview
In our previous paper [25], we performed a comparative study of three proposed algorithms for DFL under imperfect communication conditions, typified by noisy channels. These algorithms—FedNDL1, FedNDL2, and FedNDL3—differ in their handling of noise and parameter sharing, demonstrating varying degrees of resilience to communication noise. In this paper, we propose a novel algorithm that employs the Gradient Tracking method in DFL and compare its performance against the previously mentioned algorithms.
I-C Paper’s Contributions
This paper introduces a novel algorithm that employs the Gradient Tracking method in DFL, considering the impact of communication noise. Previous studies have evaluated the effectiveness of two-time scale methods in DFL with noisy channels. However, these investigations were limited by inflexible assumptions such as strong convexity in papers such as [26, 27, 28, 29]. These assumptions are rarely satisfied in practical and large-scale learning scenarios, which limits the applicability of the proposed methods. The new algorithm presented in this paper addresses these limitations by using a more flexible optimization framework that can handle smooth non-convex objective functions. We conducted experiments using the proposed algorithm in practical distributed learning scenarios, under the presence of noise. The results have shown that the algorithm is capable of reducing overall loss and mitigating consensus error, despite the presence of noise.
@@ I-C Paper’s Contributions
We evaluate the performance of decentralized FL using the Federated Model Update Tracking algorithm, which incorporates noise into tracking parameter transmissions to clients or servers. Unlike prior models where communication noise was not considered, our approach integrates noise post-local SGD updates, facilitating parameter exchange via a gossip/mixing matrix and updating the global parameters thereafter. We also compare the noise resilience of our algorithm with previously developed algorithms, providing both theoretical and empirical evidence of its robustness against communication noise [25].We show that the algorithm proposed in this paper handles the communication noise better than our previously proposed FedNDL3 algorithm.
II Problem Statement, Algorithm, and Assumptions
II-A Problem Statement
In this section, we define the framework, the underlying assumptions, and the methodologies evaluated in this study. The discussion begins with a typical DFL configuration, wherein $n$ clients possess distinct local datasets and perform consensus-based learning to obtain the global parameters. The problem is mathematically formulated as follows:
$\min_{x\in\mathbb{R}^{d}}\left[f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x_{i})\right],$ (1)
@@ II-A Problem Statement
where each $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, for $i$ ranging from 1 to $n$, signifies the local objective function for the $i^{th}$ client node. The stochastic expression of the local objective function is presented as:
$f_{i}(x_{i})=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}(x_{i},\xi_{i})],$ (2)
@@
in which $\xi_{i}$ represents the sample data drawn from the data distribution $\mathcal{D}_{i}$ specific to the $i^{th}$ client. Here, $F_{i}(x_{i},\xi_{i})$ denotes the loss function calculated for each client and their respective data sample $\xi_{i}$. The notation $x_{i}\in\mathbb{R}^{d}$ refers to the parameter vector for client $i$, while $X\in\mathbb{R}^{d\times n}$ is the matrix constituted by these parameter vectors. The fundamental goal for the clients is to collaboratively attain a state of optimality, denoted as $x_{i}=x^{*}=\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}f(x)$, which represents the global minimum.
@@
The mixing matrix, $W=[w_{ij}]\in[0,1]^{n\times n}$, is a non-negative, symmetric, that is, $W=W^{\top}$, and doubly stochastic matrix, that is, $W\mathds{1}=\mathds{1},\mathds{1}^{\top}W=\mathds{1}^{\top}$, where $\mathds{1}$ is the column vector of size $n$ whose elements are ones.
FedNDL1 [25]: In FedNDL1, the model updates are performed by each client in parallel and then each client then communicates the parameters to the neighbors. The communication is topology dependent. The neighboring client receives a noisy version of the parameters due to the imperfect communication channel,
@@
where $x_{j}^{(t+\frac{1}{2})}$ is the vector of parameters sent by client $j$ and $\delta_{j}^{(t)}\in\mathbb{R}^{d}$, is a zero mean random noise. We assume the noise to have a zero mean, the noise variance is
$D^{2}_{t,j}=\operatorname{\mathbb{E}}[\|\delta_{j}^{(t)}\|^{2}].$
FedNDL2 [25]: Similar to the previous algorithm, this algorithm also performs a two-stage process. However, In FedNDL2, the consensus step is performed before computing the individual gradients and parameters,
$x_{i}^{(t+\frac{1}{2})}=\sum_{j=1}^{n}w_{ij}(x_{j}^{(t)}+\delta_{j}^{(t)}),$ (4)
FedNDL3 [25]: In FedNDL3, the clients share their gradients over a noisy communication channel instead of the weights followed by the SGD update. This idea comes from our Noisy-FL motivation and the fact that SGD is inherently a noisy process. So, pursuing this scenario gives more flexibility to handle the noise as a part of the SGD process. The entire formulation for this algorithm can be written as,
@@
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 1: Loss versus iterations for various $\mu$ values for different communication topologies (No noise scenario).
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 2: Consensus error versus iterationsfor various $\mu$ values for different communication topologies (No noise scenario).
FedNMUT: We next describe our proposed algorithm studied in this paper. The global update tracking algorithm aims to gain the advantages of gradient tracking while eliminating communication overhead. Instead of exchanging gradients, clients communicate model updates to their neighbors.
In the Global Update Tracking (GUT) approach, the model updates, $x_{i}^{t}-x_{i}^{t-1}$, are transferred instead of the gradients $g_{i}^{t}$. Thus, this approach involves each client $i$ transmitting its model updates to its neighbors, thereby avoiding the direct communication of model gradients. Each client maintains a record of its neighbor’s model states as $\hat{x}_{j}$, updating this information with incoming model updates to maintain the latest state information, as outlined in line 10 of the algorithm.
In this algorithm, the variable $\Delta_{i}^{t}$, is calculated for each client $i$, which aggregates the local gradient update $g_{i}^{t}$ and the gossip averaging update $\sum_{j}(w_{ij}-I_{ij})\hat{x}_{j}^{t}$, as shown in line 5 of the Algorithm 2. The tracking variable $y_{i}^{t}$ is calculated as per line 6 in the algorithm, factoring in both the local gradient and the gossip averaging updates represented by $\Delta_{i}^{t}$. The gossip component of the update for each client $i$, expressed as $\sum_{j}w_{ij}(\hat{x}_{j}^{t}-x_{i}^{t})$, is based on the client’s own model parameters. To integrate this in the calculation of the tracking variable $y_{i}^{t}$, the information from neighboring clients ($y_{j}^{t}$) must be adjusted to reflect the client’s own reference frame, resulting in an added term $\frac{1}{\eta}(\hat{x}_{j}^{t}-x_{i}^{t})$ in the update formula mentioned in line 6 of the algorithm. This term can be consolidated and viewed as an bias term, $b^{t}$ as shown in equation (7) below. Additionally, to optimize the effectiveness of this method, the correction factor for the tracking variable is scaled by $\mu$, a hyperparameter that can be adjusted to maximize the algorithm’s performance.
The term denoted $\delta^{(t)}_{i}$ represents the communication noise that is added to the tracking variable $y_{i}^{t}$ shown in line 7 of the algorithm. The noisy tracking variable is then transmitted to neighboring clients as shown in lines 8–10 of the algorithm. Setting the hyperparameter $\mu$ allows us to replicate the FedNDL3 algorithm update as per equation (5).
We summarize the above described algorithms in a compact way as Algorithm algorithm 1.
Algorithm 1 FedNDL1, FedNDL2, and FedNDL3
@@
1:  Input: For each node $i$ initialize: $x_{i}^{(0)}\ \in\ \mathbb{R}^{d}$, step size $\{\eta_{t}\}_{t=0}^{T-1}$, mixing matrix $W$, noise from the communication channel $\delta^{(t)}$
2:  for $t=0,\dots,T$ do
3:      FedNDL1:
4:      Run in parallel for each client $i$
5:      Sample $\xi_{i}^{(t)}\mbox{, compute }g_{i}^{(t)}=\widetilde{\nabla}f_{i}(x_{i}^{(t)},\xi_{i}^{(t)})$
6:      $x_{i}^{(t+\frac{1}{2})}\ =\ x_{i}^{(t)}-\eta_{t}g_{i}^{(t)}$
7:      $x_{i}^{(t+1)}\ =\ \sum_{j=1}^{n}w_{ij}\ (x_{j}^{(t+\frac{1}{2})}+\delta_{j}^{(t)})$
8:      FedNDL2:
9:      $x_{i}^{(t+\frac{1}{2})}\ =\ \sum_{j=1}^{n}w_{ij}\ (x_{j}^{(t)}+\delta_{j}^{(t)})$
10:      Run in parallel for each clients $i$
11:      Sample $\xi_{i}^{(t)}\mbox{ , }g_{i}^{(t+\frac{1}{2})}=\widetilde{\nabla}f_{i}(x_{i}^{(t+\frac{1}{2})},\xi_{i}^{(t)})$
12:      $x_{i}^{(t+1)}\ =\ x_{i}^{(t+\frac{1}{2})}-\eta_{t}g_{i}^{(t+\frac{1}{2})}$
13:      FedNDL3:
14:      Run in parallel for each client $i$
15:      Sample $\xi_{i}^{(t)}\mbox{, compute }g_{i}^{(t)}=\widetilde{\nabla}f_{i}(x_{i}^{(t)},\xi_{i}^{(t)})$
16:      $x_{i}^{(t+1)}\ =\ x_{i}^{(t)}-\eta_{t}\sum_{j=1}^{n}w_{ij}\ (g_{j}^{(t)}+\delta_{j}^{(t)})$
17:  end for
The FedNMUT algorithm can be formulated as
$x_{i}^{(t+1)}=x_{i}^{(t)}-\eta_{t}\widetilde{y_{i}}^{t}=x_{i}^{(t)}-\eta_{t}({y_{i}}^{t}+\delta_{i}^{t}),\$ (6)
where
$y_{i}^{t}=\Delta_{i}^{t}+\mu\underbrace{\Big{[}\sum\limits_{j\in\mathcal{N}(i)}w_{ij}(\widetilde{y_{j}}^{t-1}-\frac{1}{\eta}(\hat{x}_{j}^{t}-x_{i}^{t}))-\Delta_{i}^{t-1}\Big{]}}_{b^{t}},\$ (7)
and
$\Delta_{i}^{t}=g_{i}^{t}-\frac{1}{\eta}\sum_{j\in\mathcal{N}(i)}w_{ij}(\hat{x}_{j}^{t}-x_{i}^{t}).$
Algorithm 2 FedNMUT - Noisy Model Update Tracking
1:  Input: For each node $i$ initialize: $x_{i}^{(0)}\ \in\ \mathbb{R}^{d}$, step size $\{\eta_{t}\}_{t=0}^{T-1}$, neighbors’ copy $\hat{x}_{j}^{0}$, step size $\eta$, scaling factor $\mu$, mixing matrix $W=[w_{ij}]_{i,j\in[1,n]}$, $\mathcal{N}(i)$ represents neighbors of $i$ including itself, noise from the communication channel $\delta^{(t)}$
@@
II-B Assumptions
The assumptions used in the convergence analysis of the proposed FedNMUT decentralized algorithm are commonly used in the literature, see [3, 2, 30].
Assumption 1 (Smoothness).
The objective function $F_{i}(x,\xi)$ is $L$-smooth with respect to $x$, for all $\xi$. Each $f_{i}(x)$ is $L$-smooth, that is,
[…]
Hence the function $f$ is also $L$-smooth.
Assumption 2 (Bounded Variance).
The variance of the stochastic gradient of each client $i$ is bounded, that is,
$\operatorname{\mathbb{E}}[||{\nabla}F_{i}(x_{i}^{t},\xi_{i}^{t})-\nabla f_{i}(x_{i}^{t})||^{2}]\leq\sigma^{2},$
where $\xi_{i}^{t}$ denotes random batch of samples in client node $i$ for $t^{th}$ round, and ${\nabla}F_{i}(x_{i}^{t},\xi_{i}^{t})$ denotes the stochastic gradient. In addition, we also assume that the stochastic gradient is unbiased, that is, $\operatorname{\mathbb{E}}[{\nabla}F_{i}(x_{i}^{t},\xi_{i}^{t})]=\nabla f_{i}(x_{i}^{t})$.
Assumption 3 (Mixing matrix).
For $\rho\in(0,1]$, the mixing matrix W satisfies,
$\|(\bar{X}-X)W\|_{F}^{2}\leq(1-\rho)\|\bar{X}-X\|_{F}^{2},$
Assumption 4 (Noise model).
The noise present due to contamination of communication channel $\delta_{i}^{(t)}$ is independent, has zero mean and bounded variance, that is,
         $\operatorname{\mathbb{E}}[\delta_{i}^{(t)}]=0$ and $\operatorname{\mathbb{E}}[||\delta_{i}^{(t)}||^{2}]=D_{t,i}^{2}<\infty$.
The above assumption is specific to the imperfect information sharing setup and is also used in [28, 29, 31, 32].
Assumption 5 (Bounded Client Dissimilarity (BCD)).
For all $x\in\mathbb{R}^{d}$,
$\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(x)-\nabla f(x)\|^{2}\leq\zeta^{2}$
for some constant $\zeta$.
@@ II-B Assumptions
Figure 3: Loss versus iterations with and without noise (Var=0.005) for different communication topologies. $\mu$ = 0.02